HUBBLE SPACE TELESCOPE DAILY REPORT #985 PERIOD COVERED: 0800 EST Daily Status Report as of 277/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 4848 (GTO Low Latitude Parallel Program - Cycle 3 High) The Wide Field Camera was used to make three high latitude medium-deep sky survey observations taken in parallel with the prime science observations. All three observations were received and displayed. They were reported to look okay. 1.2 Completed HSP/POL 4888 (Remnant Stars in Supernova Remnants - Augmentation - Cycle 3 High) The High Speed Photometer was used to make 13 ultraviolet polarmetric observations of supernova remnant SS433 which has emission lines that vary by relativistic amounts in a 162 day cycle. All 13 observations were received and displayed. They were reported to look okay. 1.3 Completed PC 5072 (WF/PC Cycle 3 Calibration: PC Flats Wide. 1) Both the Wide Field Camera and the Planetary camera were used to make four Earth Flat calibration exposures using the wide and medium band filters. Three observations were received and displayed in the OSS. Two observations were underexposed, and two were normal. 1.4 Completed FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) Fine Guidance Sensor #3 was used to make astrometric observations of the binary star system HD166988 to obtain its orbital elements. These observations are important since most of the estimates of stellar mass are made by observing elements of these systems. The astrometry was monitored and no problems were seen in the telemetry. 1.5 Completed Two Sets of FOS/RD 4941 (Geometry and Generalizability of the Reflected Light Model for Seyfert Galaxies) The Faint Object Spectrograph was used to make observations of the Seyfert Galaxy MARK 1066. The observations were monitored in the OSS, and the data quality was assessed. The observation was reported to be satisfactory. 1.6 Completed Three Sets of PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make two standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. The observations were executed as scheduled. One image was reported to be partially saturated, and the other was okay. 1.7 Completed Three Sets of WFC 4381 (HST Medium - Deep Survey - Cycle 3) The Wide Field Camera was used to make two set of medium-deep high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. 1.8 Completed Two Sets of HSP/UV2 4951 (King, Augmentation, Stellar Populations, Part 2) The High Speed Photometer (Ultraviolet Side 2) was used to make observations of globular cluster NGC-6624 to make ultraviolet observations looking for a counter part to the X-Ray burster. This binary system has an eleven minute period at X-Ray wavelengths. The observations were executed as scheduled, and no problems were reported. 1.9 Completed WFC 5075 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 1) Observations were made to make a standard calibration to obtain Earth flats using the Wide Field Camera. The observations were taken as scheduled, and no problems were reported. 1.10 Completed PC 4805 (The Ultraviolet Emission Spectra of the H II Region) The Faint Object Spectrograph (Both red and Blue Sides) were used to make two ultraviolet spectral observations fo the Hour-Glass Section (i.e., the H II emission region) of the Lagoon Nebula (aka, M-8). The intent is to determine the composition and structure of this H II region. The onboard target acquisition was successful. No problems were reported with the spectral observations. 1.11 Completed FOS/BL 3908 (Sleuthing the Dynamo - Cycle 2 Continuation) The Faint Object Spectrograph was used to make spectral observations of the very young solar type star STAR-0326+ to obtain high signal to noise ratio spectra of a star similar to, but much younger than, out Sun. Five observations were received and displayed, there was a recentering event which occurred during the last observation. The other observations were reported to look normal. 1.12 Completed WFC PC 5026 (WF/PC cycle 3 Calibration: Rapid Internal Monitor) The Wide Field Camera and the Planetary Camera were used to make internal flat calibration images. These images will be used to look for measles and daisies (i.e., different forms of contamination), and to monitor scattered light. The observations were executed as scheduled, and no problems were reported. 1.13 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.14 Completed FOC/96 5045 (FOC Observations to Monitor Desorption - Cycle 3) The Faint Object Camera was used to make observations of the standard star GSC4017-38 using filter f486N in the F/96 mode to monitor the focus of the HST optics. Nine observations were received and displayed. Two of the observations were reported to have had recentering events during the observations. 1.15 Completed FGS 5059 (Seven Points of Light - Cycle 2 Calibration - Fixed) The Fine Guidance Sensors were used to make observations of standard star UPGREEN-69. The objective was to calibrate the single star interferometeric fringe pattern at seven different points in the central region of the astrometric FGS. The astrometric observations were monitored and no problems were seen in the telemetry. 1.16 Completed GHRS/2 3472 (The Properties of Single Interstellar Clouds - Cycle 2) The High Resolution Spectrograph was used to make Echelle spectral observations of the broad-lined, early-type star HD037128. Both the Echelle and grating G160M were used. All of the observations were executed, received and displayed without problems. The target was reported to be well centered. 1.17 Completed PC 4712 (WF/PC Cycle 2 Calibration: Measles Monitor) The Planetary Camera was used to make three Earth Calibration Flat images to monitor fo a form of contamination known as measles. The observations used filter F517N, both with and without filter F122M. The OSS monitored the observation and assessed the data quality. No problems were reported. The observations were reported to look okay. 1.18 Completed GHRS/2 3859 (Nighttime Measurements of Nitric Oxide - Continuation of 2544) The High Resolution Spectrograph was used to make observations of the bright B class star HD38771. It is hoped that these observations will help learn more about the diffusion processes in the Earth's upper atmosphere by studying the ultraviolet absorption bands produced by nitric oxide along the line of sight of the target star. The observations were executed as scheduled, and no problems were reported. 1.19 Completed HSP/UV2 4502 (Quasi- Periodic Oscillations in AM Herculis Binaries - Cycle 3 Medium) The High Speed Photometer was used to make time series photometric observations of the AM Her class star QQ-Vulpecula. This star is a binary star in which material from one star is accreted unto the surface of a white dwarf companion. It can very by several magnitudes over a time scale of several minutes to months. The observations were completed as scheduled, and they all were reported to look fine. 1.20 Completed Two Sets of PC 5071 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 1) The Planetary Camera was used to make four Earth calibration flat images which will be used to remove the pixel to pixel variations in gain of the chips. The flats from this proposal will be used for flat field corrections for subsequent science data. The observations were executed as scheduled. One observations was reported to be saturated, and the other three were okay. 1.21 Completed GHRS/2 3880 (Circumstellar and Interstellar Absorption Lines in Planetary Nebula) The High Resolution Spectrograph was used to study the central star in planetary nebula NGC-40 The principal investigator is not interested in the central star, but instead she is using the starlight to probe the neutral material which surrounds the ionized shell which surrounds the central star. The OSS monitored the observations and assessed the data quality. No problems were reported. 1.22 Completed Two Sets of FGS 2565 (Linking the HIPPARCOS Satellite to the Extragalactic Reference Frame Part 1 of 6) The Fine Guidance Sensors were used to make two sets of astrometric observations of objects 0637-441IN. This object lies near the apparent position of extragalactic objects. This object will also be observed with the HIPPARCOS satellite. The intent is to tie the HIPPARCOS system reference frame to the extragalactic reference frame. The observations were executed as scheduled, and two of the observations had the Stop Flag raised during the observations. The rest of the observations were successful. 1.23 Completed Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The Wide Field Camera was used to make ten Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. All ten observations were received and displayed. No anomalies were reported. 1.24 Completed GHRS/2 4869 (Spectroscopy of Interacting Binaries - 4185, Cycle 3 High) The High Resolution spectrograph was used to make observations of the interacting binary star V-Sge. IUE spectra showed the possible presence of red-shifted emission at a velocity of about 700 Km/Sec. This suggests a major mass transfer event is in process. The target was successfully located. The wavelength calibrations and spectral observations were received and displayed. They were reported to look okay. 1.25 Completed FGS 5060 (Linking the HIPPARCOS Satellite to the Extragalactic Reference Frame Part 1 of 6) The Fine Guidance Sensor were used to make astrometric observations of the object 2128-123IN. The observations will be used to determine the rotation of the reference frame used by the HIPPARCOS astrometry satellite. No problems were seen with the astrometric observations. No anomalies were noted. 1.26 Completed FOS/RD 5002 (The Inner Regions of Quasars - Repeat for HOPR #0104) The Faint Object Spectrograph was used to make five spectral observations of the inner regions of the radio loud quasar PKS2216-03. This was a repeat of proposal 4052. All five observations were received by the science recorder playback, and displayed in the OSS. They were reported to appear normal. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 20 Successful : 20 FGS #1 failed from Fine Lock to Course Track at 276/1632z. Subsequence observations were successful. Scheduled Reacquisitions : 20 Successful : 20 2.2 FHST Updates: Scheduled : 46 Successful : 46 2.3 Significant HSTARs: None 2.4 Operations Note: FGS #2 was successfully used as a secondary FGS for observations starting a 274/0812Z. The next four subsequent acquisitions were also successful. The FGS functioned as expected and no problems were reported. This FGS will now be scheduled for other subsequent science observations in the near future. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.2 Complete HSP/UV2 5005 (Gravitational Lenes - Part II _ Cycle 2 High) 3.3 Complete HSP/UV1 1091 (Ultraviolet Pulsations From X-Ray Pulsars) 3.4 Complete FGS 5054 (High Speed Astrometry - A Search for Planetary Companions to Low mass Stars) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-